OpenStudy (anonymous):

a car salesman earns a 27,200 salary each year plus commission of 320 on each car he sells. what is the minimum number of cars he must sell each year to earn at least 56,000

OpenStudy (jim766):

if the number of cars is C and his Salary is at least >= 27,200 + 320C >= 56,000 solve for C

OpenStudy (anonymous):

Fixed + Commission $27200 + 320x (x=number of cars sold) 27200 + 320x = 56000 320x = 28800 x = 28800/320 x = 90

OpenStudy (jim766):

the question says at least......he would have to sell 90 cars to make exactly 56,000 but he sold more cars he would make more. \[x \ge90\]

OpenStudy (anonymous):

read the question carefully it says minimum cars sold to make 56000, the minimum is just 90 cars to be right on $56000.

OpenStudy (jim766):

right....good call!

OpenStudy (anonymous):

haha lol, sorry, I was just sure on this one.